i must simplify (2+3i)(2-3i) i think the answer is 4-9i is that correct
Found 3 solutions by ankor@dixie-net.com, Grinnell, richwmiller:
Answer by ankor@dixie-net.com(22740)

i must simplify (2+3i)(2-3i) i think the answer is 4-9i is that correct
:
If you FOIL this you have:
4 - 6i + 6i - 9(i^2)
:
6i's cancel and i^2 = -1 so you have:
4 - 9(-1)
:
4 + 9 = 13 is the solution

Answer by Grinnell(63)

(x+2)(x-2)=x^2-4
Let's use this as a model!
remember i^2 is NEGATIVE 1!!!!!!!
We are left with 4-9i^2
4-(-9)=13,
I think you forgot to square the i--3iTIMES(negative_3i).
Does this make sense to you?
For the previous problem I got 2-i(all)/3-4i. Got confused about your numerator,
Let me hear back from you. Put 'numerator' in your response, OK?

Answer by richwmiller(17219)

foil
f=2*2
o=-3i*2
i=2*3i
l=-9i^2=-9*(-1)=+9
4-6i+6i+9=13